write a quadratice equation having the given numbers as solutions -7 and 1
use standard form, x as a variable

Answer by stanbon(75887)

write a quadratice equation having the given numbers as solutions -7 and 1
use standard form, x as a variable
---------------
x = -7 so x+7 = 0
x = 1 so x-1 = 0
----------------------
Multiply to get:
y = (x+7)(x-1)
y = x^2+6x-7
================
Cheers,
Stan H.